Hi, I need some help, would you be able to help me with this?
Solve the following system by addition.
2x + 5y = 10
5x - 2y = 10

Answer by nabla(475)

2x + 5y = 10
5x - 2y = 10
----->Add 2/5 the first equation to the second.
2x + 5y = 10
(29/5)x = 14 implies x=70/29
Now put that value for x into the first equation.
2(70/29)+5y=10
140/29+5y=10
5y=150/29 implies y=150/145=30/29
giving solution set (70/29,30/29). Check that on your own by plugging the values into both equations.